Below is the data for the exam scores of a sample of 9 students. The population is normally distributed.

85, 70, 66, 95, 77, 94, 65, 79, 84

a)Find the margin of error for a 95% confidence interval for the mean. Round to 3 decimal places.  

b)Find the confidence interval for the mean. Round to 3 decimal places.  
Answer #1

For given sample, mean is - Στ - 79.4444

Create the following table.

data data-mean (data - mean)2
85 5.5556 30.86469136
70 -9.4444 89.19669136
66 -13.4444 180.75189136
95 15.5556 241.97669136
77 -2.4444 5.97509136
94 14.5556 211.86549136
65 -14.4444 208.64069136
79 -0.4444 0.19749136
84 4.5556 20.75349136

Find the sum of numbers in the last column to get. xi – 7)2 = 990.2222

So standard deviation is Στί - )2 V - = 11.1255 η -1

a. t table value for 8df is TINV(0.05,8)=2.306

So Margin of Error is E=t* 11.1255 = 2.306 * – -= 8.552 Vn

b. Hence CI is CI = T E = 79.444 + 8.552 = (70.892, 87.996)